OpenStudy (anonymous):

solve the eqation x+3y=45 2x+y=45

mathslover (mathslover):

solve for x in the first equation ... \[\large{x+3y=45}\] \[\large{x=45-3y}\] put this in the second equation at the place of x

OpenStudy (anonymous):

Substitution... Solve for x in the first equation: x + 3y = 45 x = -3y + 45 Substitute the value of x into the second equation to solve: 2x+y=45 2(-3y + 45) +y=45 Can you solve from there?

mathslover (mathslover):

\[\large{2x+y=45}\] \[\large{2(45-3y)+y=45}\] \[\large{2(45)-2(3y)+y=45}\] \[\large{90-6y+y=45}\] \[\large{90-5y=45}\]

mathslover (mathslover):

\[\large{5y=90-45}\]

OpenStudy (anonymous):

ummm @mathslover are you giving me the final answer that i need to do the last step on??

mathslover (mathslover):

Did i proceed to the final answer?

mathslover (mathslover):

what do you want? : a) answer b) solution ... (explained) ? c) none

OpenStudy (anonymous):

a and b

mathslover (mathslover):

ok so can you solve : 90 - 5y = 45 ????????

OpenStudy (anonymous):

yeah

mathslover (mathslover):

what's the answer then ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

is that right??? @mathslover

OpenStudy (anonymous):

@jazy am i right?

OpenStudy (anonymous):

Right! (:

OpenStudy (anonymous):

okay then what do i do next?

OpenStudy (anonymous):

Well you just solved for y. y = 9 You need to find the value of x. To do that, just put the y-value into any one of the equations and solve for x. \[x+3y=45 \]\[x+3(9)=45\]\[x+27=45\]\[Solve for X\]

OpenStudy (anonymous):

so i have to subtract 27 on both sides.. when i did that i got 18

OpenStudy (anonymous):

Good (: So now you know that x = 18 y = 9

OpenStudy (anonymous):

ohhhhh okay thanks :))

OpenStudy (anonymous):

welcome. (: